OpenStudy (anonymous):

What is the product of the square root of 6(8b the square root of 36-4b) Simplify

OpenStudy (anonymous):

Rewritten to: \[\sqrt{6}(8b \sqrt{36}-4b)\]

OpenStudy (anonymous):

Okay I worked it out I think: \[\sqrt{6}(48b-4b) \]

OpenStudy (anonymous):

\[\sqrt{6}(44b)\]

OpenStudy (anonymous):

\[44b \sqrt{6} ??\]

OpenStudy (imstuck):

You're right! And here I was typing up this huge explanation and you got it all on your own! Yay!

OpenStudy (anonymous):

Ok kinda proud of myself now!